A ball is thrown vertically download with velocity 5 metre per second from the top of the tower with what velocity should another ball be thrown vertically down after 2s from the top of the same tower so that it can hit the first ball in further 2 s .

Grade:11

3 Answers

Vikas TU
7 years ago
at 2s for first ball,
h = 5*2 + 0.5g*4 = > 30 meter.
and v = 5 + 10*2 = > 25 m/s
after further 2 seconds,
x= 25*2 + 0.5*g*4 = > 70 meter.
total distance it covered, => 70 + 30 => 100 meter.
for second ball,
100 = u*4 + 0.5g*16
4u = 100 – 80 = 20
4u = 20
u = 5 m/s
Prabhav Sharma
7 years ago
but answer given is 40m /s.
(1)40m/s
(2)55m/s
(3)15m/s
(4)25m/s
please check answer .thanks
----------------------------------------------------------------------------------------------------------------------------------------------------------------
Vikas TU
7 years ago
Edit:
the second ball needs to cover that 100 m distance in 2 seconds not in 4.
Therefore,
100 = u*2 + 0.5*g*4
u = 40 m/s